OpenStudy (clamin):

What is the interquartile range of the following data set? 45, 12, 48, 96, 61, 84, 29, 1, 72, 5, 14 A. 33 B. 27 C. 45 D. 60

OpenStudy (anonymous):

Okay. Check this out! It should help! http://www.statisticshowto.com/interquartile-range/ @clamin

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @clamin

OpenStudy (clamin):

I keep getting 24 but its not in the choices

OpenStudy (texaschic101):

put the numbers in order... 1,5,12,14,29,45,48,61,72,84,96 Q2 -- the middle number = 45 Q1 = 12 Q3 = 72 interquartile range is Q3 - Q1 = 72 - 12 = 60